OpenStudy (anonymous):

A man weighs 750 N on the surface of the earth. What would be his weight when standing on the moon? The masses of the earth and the moon are respectively, 5:98×1024 kg and 7:36×1022 kg. Their radii are respectively 6:37×103 km and 1:74×103 km

OpenStudy (anonymous):

Fg=GmM/R^2 Fg(Earth)=GmM(Earth)/R(Earth)^2; Fg(moon)=GmM(moon)/R(moon)^2 =>Fg(m)/Fg(E)=M(m)*R(E)^2/(M(E)*R(m)^2)=0,16495 => Fg(Moon)=0,16495*750=123,7N

OpenStudy (anonymous):

thanks so so much